HIGH-ORDER TERMS IN THE ASYMPTOTIC EXPANSIONS OF

THE STEADY-STATE VOLTAGE POTENTIALS IN THE PRESENCE OF CONDUCTIVITY INHOMOGENEITIES OF SMALL

DIAMETER

HABIB AMMARI HYEONBAE KANG

Abstract. We derive high-order terms in the asymptotic expansions of the steadystate voltage potentials in the presence of a finite number of diametrically small inhomogeneities with conductivities different from the background conductivity. Our derivation is rigorous, and based on layer potential techniques. The asymptotic expansions in this paper are valid for inhomogeneities with Lipschitz boundaries and those with extreme conductivities.